North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotives Q6 63395 (NER T2 2238) Q6 Blog Q6 - 15/16 February 2022

Q6 - 15/16 February 2022

Ian reports (and I can confirm), that Tuesday was a cold, wet, morning at Grosmont, but it became brighter before lunchtime and was a lovely sunny afternoon. Present were Steve Hyman, Chris Lawson, Gordon Wells and Ian Pearson. Bill Dobson popped in at lunch time to check on how things were going and have his lunch.

Gordon worked on the Q6, altering oil pipe work for the new atomisers that are being fitted. This entailed cutting the existing pipes, shortening them, and brazing new fittings. These pipes had been annealed last week, and were blown through with compressed air by Ian on Tuesday morning. All 4 pipes have now been temporarily fitted to the atomisers on the new bracket. Note that these 4 pipes feed from the atomisers to the valves. The pipes from the lubricator to the atomisers will be sorted next week. Steve completed greasing the Q6 tender. No problems today thanks to the tip from Ian Storey to use a welding rod to release air out of the grease gun. Thanks Ian - it really does work.

Wednesday was a mild, stormy, day with strong winds and heavy showers as the start of storm Dudley. The side door had to be tied to the seat outside to stop it banging shut. The team of working enthusiasts were Bill Dobson, Trevor Wilford, Nigel Hall and Ian Pearson.

Ian's first job was to put all our lifting equipment out for its six-monthly inspection, which included strops, shackles, pull lifts, small jack, lifting table, transmission jack and 3 ton chain block. Every item passed OK and are now colour coded Green. Trevor, liaised with Bill on the items that he was required to make up for the drop grate. Various items were made ready for welding and fitting inside the ashpan area to enable the grate to move correctly. While Bill managed to get inside the firebox and gain access into the ashpan to remove the main operating rod and fittings, these need alteration and new fittings to make the grate operate correctly.

A box of atomiser fittings was delivered to the MPD from Barry Gamble's steam supplies. Bill ordered these items only last week. He collected the items from the office and brought them to Deviation workshop and checked the order was correct. Ian removed the two-atomiser units that Gordon had fitted on Tuesday. These required four new atomiser parts that had just been delivered, called swirls or baffles. They allow the oil to mix with steam and spray the mixture into the valves. These are to be fitted next week.